OpenStudy (chycora):

Carmelo bought an electric guitar for $240.00 during the yearly sale at Strings Music Store. He purchased the guitar for $75.00 less than one third of its original price. Choose the equation you would use to find the original price, x, of the guitar.

OpenStudy (chycora):

@jackmullen55 @Jaynator495

jaynator495 (jaynator495):

Hulk Doesnt Do Math x_x

OpenStudy (chycora):

awww

OpenStudy (chycora):

:( ok

OpenStudy (chycora):

@Michele_Laino

jaynator495 (jaynator495):

My question... how in the heck are you 70ss over night? LOL

OpenStudy (michele_laino):

Hint: If I call with x the original price of the guitar, then one third of the original price is: x/3 and $75 less than one third of the original price is : x/3 - 75

jaynator495 (jaynator495):

Just yesterday i was barfing rainbows on you being your new... next thing i know your already grown up LOL

OpenStudy (chycora):

the answer is C right @Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (chycora):

can you help me with more and thanx

OpenStudy (michele_laino):

ok!

OpenStudy (chycora):

thanx

OpenStudy (chycora):

Factor completely. 42 + 51x - 18x² A. 6(3x + 1)(7 - x) B. (3x - 7)(2x - 2) C. 3(2 + 3x)(7 - 2x) D. 3(3x - 7)(2x - 2)

OpenStudy (michele_laino):

here we have to find the roots of the polynomial: \[18{x^2} - 51x - 42 = 0\]

OpenStudy (chycora):

ok

OpenStudy (chycora):

18^2=324

OpenStudy (michele_laino):

no, we have to apply this formula: \[\Large x = \frac{{ - b \pm \sqrt {{b^2} - 4 \times a \times c} }}{{2 \times a}}\] where a= 18, b=-51, and c= -42

OpenStudy (chycora):

\[x=\frac{ -51\pm \sqrt{-51^{2}- 4\times18\times-42} }{ 2*18 }\]

OpenStudy (chycora):

is that right

OpenStudy (michele_laino):

yes! \[x = \frac{{51 \pm \sqrt {{{51}^2} - \left( {4 \times 18 \times - 42} \right)} }}{{2 \times 18}}\]

OpenStudy (michele_laino):

now, we have: \[\Large \begin{gathered} x = \frac{{51 \pm \sqrt {{{51}^2} - \left( {4 \times 18 \times - 42} \right)} }}{{2 \times 18}} = \hfill \\ \hfill \\ = \frac{{51 \pm \sqrt {2601 + 3024} }}{{36}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (chycora):

\[=\frac{51\pm5625 }{ 36 }\]

OpenStudy (michele_laino):

I got this: \[\begin{gathered} \frac{{51 \pm \sqrt {2601 + 3024} }}{{36}} = \frac{{51 \pm \sqrt {5625} }}{{36}} = \hfill \\ \hfill \\ = \frac{{51 \pm 75}}{{36}}...? \hfill \\ \end{gathered} \]

OpenStudy (chycora):

ok so wat kn

OpenStudy (michele_laino):

we have the subsequent roots: \[\Large \begin{gathered} {x_1} = \frac{{51 + 75}}{{36}} = \frac{{126}}{{36}} = \frac{7}{2} \hfill \\ \hfill \\ {x_2} = \frac{{51 - 75}}{{36}} = - \frac{{24}}{{36}} = - \frac{2}{3} \hfill \\ \end{gathered} \]

jaynator495 (jaynator495):

I came back and realized it was a different person who started yesterday... my bad x'D

OpenStudy (chycora):

is the answer c @Michele_Laino cuz i think so

OpenStudy (michele_laino):

in ordert o find the answer you have to do this computation: \[\Large \begin{gathered} - 18\left( {x - \frac{7}{2}} \right) \times \left( {x + - \frac{2}{3}} \right) = \hfill \\ \hfill \\ = - 3 \times 2\left( {x - \frac{7}{2}} \right) \times 3 \times \left( {x + \frac{2}{3}} \right) = ... \hfill \\ \end{gathered} \]

OpenStudy (chycora):

o shot i took to long on this question and it loged me out so nvm

OpenStudy (michele_laino):

Sorry, I have made a typo, here is the right step: \[\Large \begin{gathered} - 18\left( {x - \frac{7}{2}} \right) \times \left( {x + \frac{2}{3}} \right) = \hfill \\ \hfill \\ = - 3 \times 2\left( {x - \frac{7}{2}} \right) \times 3 \times \left( {x + \frac{2}{3}} \right) = ... \hfill \\ \end{gathered} \]

OpenStudy (chycora):

ur ok

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

hint: you have to multiply the factor 2 by the first parentheses, and then you have to multiply the factor 3 by the second parentheses

OpenStudy (chycora):

ok thanx ur a big help

OpenStudy (michele_laino):

you should get this: \[\Large \begin{gathered} - 18\left( {x - \frac{7}{2}} \right) \times \left( {x + \frac{2}{3}} \right) = \hfill \\ \hfill \\ = - 3 \times 2\left( {x - \frac{7}{2}} \right) \times 3 \times \left( {x + \frac{2}{3}} \right) = \hfill \\ \hfill \\ = - 3 \times \left( {2x - 7} \right) \times \left( {3x + 2} \right)... \hfill \\ \end{gathered} \]